OpenStudy (anonymous):

the amout P of pollution varies directly with the population N of people. City A has a population of 494,000 and produces 260,000 tons of polluntants. Find how many tons of pollution we should expect City B to produce if we know that the polulation is 344,000

OpenStudy (anonymous):

B/ 344,000 = 260,000 /494,000 -> B = ...